OpenStudy (anonymous):

the weight of a body above sea level varies inversely with the square of the distance from the center of earth. if a womeh weights 120 pounds when she is at sea leve, 3960 miles from the center of earth, how much will she weigh when she is at the top of a mountain 1.7 miles above sea level?

OpenStudy (anonymous):

Still 120, more or less.... To get to the actual computation of this problem,\[W(r_1)=W_0\frac{r_0^2}{r_1^2}=120 \frac{3960^2}{3961.7^2}\approx 119.9\]